Suppose we are concerned about the variability of the number of raisins in a box of raisin bran. Assume that the number is distributed normally. If a random sample of 16 boxes is taken, what is the probability that the sample variance will be more than twice the population variance?
Answer #1

here as we know that chi square test statistic X2 =(n-1)(s/σ)2   with (n-1) degree of freedom

therefore P(s2/ σ2 >2 )=P((n-1)(s/σ)2 >2*(16-1))=P(X2 >30) =0.0119